# From High-SNR Radar Signal to ECG: A Transfer Learning Model with Cardio-Focusing Algorithm for Scenarios with Limited Data

### v1

@@ Abstract
Electrocardiogram (ECG), as a crucial find-grained cardiac feature, has been successfully recovered from radar signals in the literature, but the performance heavily relies on the high-quality radar signal and numerous radar-ECG pairs for training, restricting the applications in new scenarios due to data scarcity. Therefore, this work will focus on radar-based ECG recovery in new scenarios with limited data and propose a cardio-focusing and -tracking (CFT) algorithm to precisely track the cardiac location to ensure an efficient acquisition of high-quality radar signals. Furthermore, a transfer learning model (RFcardi) is proposed to extract cardio-related information from the radar signal without ECG ground truth based on the intrinsic sparsity of cardiac features, and only a few synchronous radar-ECG pairs are required to fine-tune the pre-trained model for the ECG recovery. The experimental results reveal that the proposed CFT can dynamically identify the cardiac location, and the RFcardi model can effectively generate faithful ECG recoveries after using a small number of radar-ECG pairs for training. The code and dataset are available after the publication.
Index Terms:
Contactless Vital Sign Monitoring, Radar-Based Sensing, Transfer Learning, Derivative-Free optimization
@@ I Introduction
Radio detection and ranging (radar) system is originally designed for military detection of large aircraft by emitting electromagnetic waves and evaluating the reflections. The follow-up research has investigated the civilian use of radar systems for contactless sensing in various scenarios, such as autonomous driving [1] and human monitoring [2]. Over the past decade, radar-based sensing has been empowered by deep neural networks to process non-stationary reflected signals or high-dimensional data, enabling versatile applications to replace contact- or visual-based measurement for convenience or privacy concerns (e.g., vital sign monitoring [3], gesture recognition [4], person identification [5]).
Radar-based vital sign monitoring, as a popular branch of radar-based sensing, has been explored for decades to measure heart rate or respiration rate in a contactless manner [2], and some further studies leverage the deep neural network to realize domain transformation from cardiac mechanical activities (i.e., heartbeat) to electrical activities (i.e., electrocardiogram (ECG)), providing a fine-grained cardiac measurement for wellness monitoring or clinical diagnosis [6, 7, 8, 9, 10, 11, 12]. In the literature, radar-based ECG recovery is only realized by deep-learning-based methods, because the domain transformation is extremely complex to be modeled mathematically while such transformation can be learned by deep learning model due to the great nonlinear mapping ability [9].
Similar to other research fields involved with deep learning, radar-based ECG recovery also asks for numerous radar signals to train the deep learning model with synchronous ECG ground truths [11, 8, 13]. According to previous research, the performance of the deep learning model degrades heavily after reducing $30\%$ of the training data even after applying proper data augmentation techniques [10], causing difficulties for the deployment in new scenarios due to the demand for hours of ECG collection [7]. However, the method for reducing dependence on data quantity is rarely investigated for radar-based ECG recovery, and all the deep-learning-based ECG recovery models are trained in a supervised manner with large dataset containing $3-32$ hours of synchronous radar-ECG pairs [9, 14, 11].
In contrast, most studies are dedicated to inventing advanced signal processing algorithms to enhance the signal quality, because the deep learning model for ECG recovery is vulnerable to the inputs contaminated by noises and requires high signal-to-noise ratio (SNR) radar signals as inputs [9, 3]. The methods for capturing high-SNR radar signals can be categorized into two groups:
@@
The first type of method is not suitable for some commonly used frequency-modulated continuous-wave (FMCW) radar platforms (e.g., TI AWR-x radar) due to the on-board digital front-end module filtering the frequency-modulated feature of baseband signal (i.e., circular IQ plot) [21], preventing the broad applications of this approach in commercial radar. The second type of method relies on accurate localization of the chest region, while the existing methods only provide a rough location of the human body, causing a deviation of several decimeters due to different postures of the subject [22]. Therefore, the methods based on signal accumulation may fail because only a minority of range bins contain cardiac features, hence not subjecting to the law of large numbers [3]. Although some aforementioned studies have proposed methods for selecting or clustering the useful range bins with cardiac features [9, 11], the computational cost for traversing a large objective space can be huge without an accurate cardiac location.
@@
  • The proposed CFT algorithm has been validated on sitting subjects in various scenarios and could provide radar measurements with better SNR compared with existing methods. In addition, the pre-trained RFcardi framework can be easily adapted to realize radar-based ECG recovery with a small amount of synchronous radar-ECG measurements for fine-tuning.
@@ II-A FMCW Radar Foundations
  • Virtual antenna array (channel) is a commonly used concept in multiple-input and multiple-out (MIMO) radar systems and is able to realize complex modulations or beamforming [16]. However, this study mainly leverages the phase difference across antenna channels to estimate the AoA of the objects.
@@ II-B Cardiac Signal Extraction from FMCW Radar
At last, some common noises, such as respiration and thermal noise, can be easily removed using a band-pass filter and differentiator to make sure that the final $h(E,t)$ mostly contains cardiac-related features from point $E$ as shown in Figure LABEL:sub@fig:radar_good.
Challenge: It is natural to think the high-SNR radar signal can be searched in a constrained space by optimization, while there is no appropriate method to assess the signal SNR in terms of cardiac features contained, and the objective space is actually highly discontinuous with adjacent points may revealing totally different SNR as shown in Figure LABEL:sub@fig:radar_good and LABEL:sub@fig:radar_bad, restricting the application of common gradient-based optimization algorithms.
Figure 1: Challenges for radar-based ECG recovery: (a) and (b) Radar signals with high and low SNR for adjacent points with a distance of $0.03$m; (c) and (d) Inference results of a well-trained deep learning model in original and new scenarios.
@@ II-C Transfer Learning
In addition to high-SNR radar inputs, accurate ECG recovery also relies on the scale of datasets to train the deep learning model, with the previous research normally adopting the dataset with the length of $3-32$ hours [9, 14, 11]. Unfortunately, the initial experimental results show that the well-trained deep learning model cannot be directly used for the signal collected from new scenarios even using a similar radar configuration as shown in Figure LABEL:sub@fig:org_scen and  LABEL:sub@fig:cardi_good, because different in-door scenarios and radar configurations may have unknown interference on radar signals [2].
Inspired by other signal-based research [20, 26], transfer learning is a promising paradigm to learn the latent representation from unlabeled radar signal to capture basic cardiac features in an SSL manner, reducing the requirement of cumbersome ECG collection. Then, only a small amount of synchronous radar-ECG pairs is required to fine-tune the pre-trained model to realize the ECG recovery for new scenarios.
@@ III-A Overview of CFT-RFcardi Framework
  • The rough location acts as the initial state for the CFT algorithm, and the points within a constrained space will be evaluated to find the red CF point with best SNR as shown in Figure 2(b).
  • Signals extracted from the ten best points will be converted into spectrograms to pre-train the backbone with SSR as the pre-text task. Then, the same backbone will used for the fine-tuning stage so that the latent representations learned in the pre-trained model can be seemly transferred for the ECG recovery task, as shown in Figure 2(c).
In addition, TI-AWR 1843 radar operated at $77$Ghz with $2$ Tx and $4$ Rx will be used for data collection, enabling $8$ virtual antenna channels for high-quality signal extraction. The detailed scenario descriptions and radar configurations will be provided in Section IV-A.
@@ III-B Rough Localization
The received radar signal is first formatted as a standard data matrix in terms of different chirps, frames and virtual channels to provide measurement of range, velocity and AoA, respectively. For the current research level of radar-based vital sign monitoring, the subjects are all quasi-static without velocity, and only the range-angle (RA) map will be calculated using fast Fourier transform (FFT) as shown in Figure 2(a), with a detailed illustration of signal waveform and processing shown in Figure 3.
@@ III-B2 Angle FFT
The ability of AoA detection relies on the MIMO system using time division multiplexing (TDM-MIMO), with multiple Tx alternately transmitting chirp signals and the corresponding reflections can be distinguished during receiving as shown in Figure 3(b). Due to the physical distance varies for different Tx/Rx combinations (i.e., Tx2222Rx4444 creates $8$ virtual channels), an extra propagation $\Delta\phi_{v}$ delay will be introduced as:
@@ III-C1 Template Design for Assessing SNR
An explicit SNR can be calculated with the known “clean” signal, while the ”clean” signal for vital signs normally reveals two prominent vibrations corresponding to the ventricular contraction and relaxation [7], as shown in Figure LABEL:sub@fig:target_org. However, considering the vibrations may have subtle differences due to different scenarios or radar configurations (e.g., noise figure and sampling frequency), a universal template $h_{m}$ is designed in this study to fit the envelope of the radar signal as:
@@
with $a_{1}$, $a_{2}$ controlling the amplitudes of the peaks, $b_{1}$, $b_{2}$ determining the centers of the peaks and $c_{1}$, $c_{2}$ adjusting the width of the peaks. In practice, $a_{1}$ and $b_{1}$ will be fixed based on the dominant peaks detected as the red points in Figure LABEL:sub@fig:target_sig, and other parameters are left to be determined as a simple curve fitting problem. Finally, the mean square error (MSE) between the radar signal envelope and the synthetic template is reckoned to be an assessment of signal SNR as shown in Figure LABEL:sub@fig:target_sig, because fewer components could fit the designed template for low-SNR radar signal without obvious cardiac features, as shown in Figure LABEL:sub@fig:target_org_bad and LABEL:sub@fig:target_sig_bad.
@@ III-C2 Derivative-free Optimization (DFO)
where $\gamma_{k}>0$ is the grid size parameter and $D$ contains several vectors $p$ for possible searching directions, as shown in Figure LABEL:sub@fig:cft_1. The local convergence of CS is ensured by dense search directions $D$ and a refined grid size $\gamma_{k}$ to find better $E_{i}$ compared with current $E_{k}$ [27]. However, the highly discontinuous objective space for radar-monitored vital signs may have numerous local minima that distract the optimization algorithm, i.e., the signal SNR of the adjacent points might be very different, as shown in Figure LABEL:sub@fig:radar_good and LABEL:sub@fig:radar_bad.
@@
$S_{k}:=\{E\in G_{k}:\|E-E_{k}\|_{\infty}\leq\Gamma_{k}a\}$ (8)
with $a=\max\left\{\left\|a^{\prime}\right\|_{\infty}:a^{\prime}\in D\right\}$ and $\Gamma_{k}$ as the size parameter for the search region. An intuitive interpretation of $S_{k}$ is the point set that contains grid points inside and on the boundary of the bold line controlled by $\Gamma_{k}$, as shown in Figure LABEL:sub@fig:cft_2.
@@
  • If a new incumbent point $E_{i}$ is found with better SNR, the search region will be doubled as $\Gamma_{k+1}=2\Gamma_{k}$ (line 9), and the grid size will be empirically set as $\gamma_{k+1}=\min(\Gamma_{k},{\Gamma_{k}}^{2})$ (line 17), enabling to search in a broader space in the next iteration.
  • If there is no better point than the current $E_{k}$ on the current grids $G_{k}$, another searching stage will be performed only within the search region $S_{k}$ (line 10). Then, If a better point $E_{i}$ is found, $\Gamma_{k+1}$ and $\gamma_{k+1}$ is obtained as above (line 12), otherwise, the search region will be halved as $\Gamma_{k+1}=\Gamma_{k}/2$ (line 15) for a finer search with $\gamma_{k+1}=\min(\Gamma_{k},{\Gamma_{k}}^{2})$ (line 17).
@@
4:$E_{b}=\underset{E\in\mathbb{R}^{n}}{\arg\min}\left\{\mathcal{F}(E):E\in\Omega\right\}$
@@ III-D2 Pre-text Task Training and Fine-tuning
The pre-text task used for SSL should reveal certain inherent features in the radar-monitored vital sign, and two major features used for traditional heart rate estimation are periodicity and sparsity [2]. In this work, the duration of each segment is $4$ sec and may not reveal strong periodicity. Therefore, SSR will be used as the pre-text task in RFcardi and is defined as:
@@
where $h$ is the high-SNR radar signal, $\Phi$ means the observation matrix, $x$ is the sparse representation for heartbeats and $n$ represents the residual noise. The traditional SSR task can be seemly converted to a system identification problem by viewing $\Phi$ as a multi-channel adaptive filter, and the estimation of the filter coefficient is the same as training a CNN-based neural network (i.e., training CNN kernels) [30].
@@
$\mathcal{L}=\|x-x^{\prime}\|_{2}+\underbrace{\lambda_{s}\frac{\|x\|_{1}/\|x\|_{2}-1}{\sqrt{m}-1}}_{\text{sparse penalty}}$ (10)
where $m$ is the length of the signal, $x$ is the output the from deep learning model, $x^{\prime}$ is the sparse ground truth with values for the radar peaks maintained (1st vibration in Figure LABEL:sub@fig:target_org) while other values are set to 00. The sparse penalty has a range of $[0,\lambda_{s})$ with a smaller value indicating better sparsity [31].

### v2

@@ Abstract
Electrocardiogram (ECG), as a crucial fine-grained cardiac feature, has been successfully recovered from radar signals in the literature, but the performance heavily relies on the high-quality radar signal and numerous radar-ECG pairs for training, restricting the applications in new scenarios due to data scarcity. Therefore, this work focuses on radar-based ECG recovery in new scenarios with limited data and proposes a cardio-focusing and -tracking (CFT) algorithm to precisely track the cardiac location to ensure an efficient acquisition of high-quality radar signals. Furthermore, a transfer learning model (RFcardi) is proposed to extract cardio-related information from the radar signal without ECG ground truth based on the intrinsic sparsity of cardiac features, and only a few synchronous radar-ECG pairs are required to fine-tune the pre-trained model for ECG recovery. The experimental results reveal that the proposed CFT can dynamically identify the cardiac location, and the RFcardi model can effectively generate faithful ECG recoveries after using a small number of radar-ECG pairs for training. The code and dataset will be made available after publication.
@@ I Introduction
Radio detection and ranging (radar) system was originally designed for military detection of large aircraft by emitting electromagnetic waves and evaluating the reflections. The follow-up research has investigated the civilian use of radar systems for contactless sensing in various scenarios, such as autonomous driving [1] and human monitoring [2]. Over the past decade, radar-based sensing has been empowered by deep neural networks to process non-stationary reflected signals or high-dimensional data, enabling versatile applications to replace contact- or visual-based measurement for convenience or privacy concerns (e.g., vital sign monitoring [3], gesture recognition [4], person identification [5]).
Radar-based vital sign monitoring, as a popular branch of radar-based sensing, has been explored for decades to measure heart rate or respiration rate in a contactless manner [2], and some further studies leverage the deep neural network to realize domain transformation from cardiac mechanical activities (i.e., heartbeat) to electrical activities (i.e., electrocardiogram (ECG)), providing a fine-grained cardiac measurement for wellness monitoring or clinical diagnosis [6, 7, 8, 9, 10, 11, 12]. In the literature, radar-based ECG recovery is only realized by deep-learning-based methods, because the domain transformation is extremely complex to be modeled mathematically while such transformation can be learned by a deep learning model due to the great nonlinear mapping ability [9].
Similar to other research fields involved with deep learning, radar-based ECG recovery also asks for numerous radar signals to train the deep learning model with synchronous ECG ground truths [11, 8, 13]. According to previous research, the performance of the deep learning model degrades heavily after reducing $30\%$ of the training data even after applying proper data augmentation techniques [10], causing difficulties for the deployment in new scenarios due to the demand for hours of ECG collection [7]. However, the method for reducing dependence on data quantity is rarely investigated for radar-based ECG recovery, and all the deep-learning-based ECG recovery models are trained in a supervised manner with large dataset containing 3323-32 hours of synchronous radar-ECG pairs [9, 14, 11].
In contrast, most studies are dedicated to inventing advanced signal processing algorithms to enhance the signal quality, because the deep learning model for ECG recovery is vulnerable to the inputs contaminated by noise and requires high signal-to-noise ratio (SNR) radar signals as inputs [9, 3]. The methods for capturing high-SNR radar signals can be categorized into two groups:
@@
The first type of method is not suitable for some commonly used frequency-modulated continuous-wave (FMCW) radar platforms (e.g., TI AWR-x radar) due to the on-board digital front-end module filtering the frequency-modulated feature of baseband signal (i.e., circular IQ plot) [21], preventing the broad applications of this approach in commercial radar. The second type of method relies on accurate localization of the chest region, while existing methods only provide a rough location of the human body, causing a deviation of several decimeters due to different postures of the subject [22]. Therefore, the methods based on signal accumulation may fail because only a minority of range bins contain cardiac features, hence not subjecting to the law of large numbers [3]. Although some aforementioned studies have proposed methods for selecting or clustering the useful range bins with cardiac features [9, 11], the computational cost for traversing a large objective space can be huge without an accurate cardiac location.
@@
  • The proposed CFT algorithm has been validated on sitting subjects in various scenarios and could provide radar measurements with better SNR compared with existing methods. In addition, the pre-trained RFcardi framework can be easily adapted to realize radar-based ECG recovery with a small number of synchronous radar-ECG measurements for fine-tuning.
@@ II-A FMCW Radar Foundations
  • Virtual antenna array (channel) is a commonly used concept in multiple-input and multiple-output (MIMO) radar systems and is able to realize complex modulations or beamforming [16]. However, this study mainly leverages the phase difference across antenna channels to estimate the AoA of the objects.
@@ II-B Cardiac Signal Extraction from FMCW Radar
At last, some common noises, such as respiration and thermal noise, can be easily removed using a band-pass filter and differentiator to make sure that the final $h(E,t)$ mostly contains cardiac-related features from point $E$ as shown in Figure LABEL:sub@fig:radar_good, with prominent vibration corresponding to the R peaks in the ECG signal.
Challenge: It is natural to think the high-SNR radar signal can be searched in a constrained space by optimization, while there is no appropriate method to assess the signal SNR in terms of cardiac features contained. In addition, the objective space is highly discontinuous, and the signals extracted from adjacent points reveal totally different SNR, as indicated by the poor radar collection with much noise that buries the cardiac features (peaks) in Figure LABEL:sub@fig:radar_bad, restricting the application of common gradient-based optimization algorithms.
Figure 1: Challenges for radar-based ECG recovery: (a) and (b) Radar signals with high and low SNR for adjacent points with a distance of 0.030.03m; (c) and (d) Inference results of a well-trained deep learning model in original and new scenarios.
@@ II-C Transfer Learning
In addition to high-SNR radar inputs, accurate ECG recovery also relies on the scale of datasets to train the deep learning model, with the previous research normally adopting the dataset with the length of 3323-32 hours [9, 14, 11]. Unfortunately, the initial experimental results show that the well-trained deep learning model cannot be directly used for the signal collected from new scenarios even using a similar radar configuration as shown in Figure LABEL:sub@fig:org_scen and  LABEL:sub@fig:new_scen, because different indoor scenarios and radar configurations may have unknown interference on radar signals [2].
Inspired by other signal-based research [20, 26], transfer learning is a promising paradigm to learn the latent representation from unlabeled radar signal to capture basic cardiac features in an SSL manner, reducing the need for cumbersome ECG collection. Then, only a small number of synchronous radar-ECG pairs is required to fine-tune the pre-trained model to realize the ECG recovery for new scenarios.
@@ III-A Overview of CFT-RFcardi Framework
  • The rough location acts as the initial state for the CFT algorithm, and the points within a constrained space will be evaluated to find the red CF point with the best SNR as shown in Figure 2(b).
  • Signals extracted from the ten best points will be converted into spectrograms to pre-train the backbone with SSR as the pre-text task. Then, the same backbone will be used for the fine-tuning stage so that the latent representations learned in the pre-trained model can be seamlessly transferred for the ECG recovery task, as shown in Figure 2(c).
In addition, TI-AWR 1843 radar operated at 7777GHz with $2$ Tx and $4$ Rx will be used for data collection, enabling $8$ virtual antenna channels for high-quality signal extraction. The detailed scenario descriptions and radar configurations will be provided in Section IV-A.
@@ III-B Rough Localization
The received radar signal is first formatted as a standard data matrix in terms of different chirps, frames and virtual channels to provide measurements of range, velocity and AoA, respectively. For the current research level of radar-based vital sign monitoring, the subjects are all quasi-static without velocity, and only the range-angle (RA) map will be calculated using fast Fourier transform (FFT) as shown in Figure 2(a), with a detailed illustration of signal waveform and processing shown in Figure 3.
@@ III-B2 Angle FFT
The ability of AoA detection relies on the MIMO system using time division multiplexing (TDM-MIMO), with multiple Tx alternately transmitting chirp signals and the corresponding reflections can be distinguished during receiving as shown in Figure 3(b). Due to the physical distance varies for different Tx/Rx combinations (i.e., Tx22Rx44 creates $8$ virtual channels), an extra propagation $\Delta\phi_{v}$ delay will be introduced as:
@@ III-C1 Template Design for Assessing SNR
An explicit SNR can be calculated with the known “clean” signal, while the “clean” signal for vital signs normally reveals two prominent vibrations corresponding to the ventricular contraction and relaxation [7], as shown in Figure LABEL:sub@fig:target_org. However, considering the vibrations may have subtle differences due to different scenarios or radar configurations (e.g., noise figure and sampling frequency), a universal template $h_{m}$ is designed in this study to fit the envelope of the radar signal as:
@@
with $a_{1}$, $a_{2}$ controlling the amplitudes of the peaks, $b_{1}$, $b_{2}$ determining the centers of the peaks and $c_{1}$, $c_{2}$ adjusting the width of the peaks. In practice, $a_{1}$ and $b_{1}$ will be fixed based on the amplitude and position of the dominant peaks detected as the red points in Figure LABEL:sub@fig:target_sig, and other parameters are left to be determined as a simple curve fitting problem as an optimization problem. Finally, the mean square error (MSE) between the radar signal envelope and the synthetic template is reckoned to be an assessment of signal SNR as shown in Figure LABEL:sub@fig:target_sig, because fewer components could fit the designed template for low-SNR radar signal without obvious cardiac features, as shown in Figure LABEL:sub@fig:target_org_bad and LABEL:sub@fig:target_sig_bad.
@@ III-C2 Derivative-free Optimization (DFO)
where $\gamma_{k}>0$ is the grid size parameter and $D$ contains several vectors $p$ for possible searching directions, as shown in Figure LABEL:sub@fig:cft_1. The local convergence of CS is ensured by dense search directions $D$ and a refined grid size $\gamma_{k}$ to find better $E_{i}$ compared with current $E_{k}$ [27]. As a basis for all CS-based DFO methods, $D$ can be formed by generating the orthogonal unit vectors based on the seed vector selected from the Halton sequence to ensure an evenly distributed search direction around the $E_{k}$. However, the highly discontinuous objective space for radar-monitored vital signs may have numerous local minima that distract the optimization algorithm, i.e., the signal SNR of the adjacent points might be very different, as shown in Figure LABEL:sub@fig:radar_good and LABEL:sub@fig:radar_bad.
@@
$S_{k}:=\{E\in G_{k}:\|E-E_{k}\|_{\infty}\leq\Gamma_{k}\}$ (8)
with $\Gamma_{k}$ as the size parameter for the search region. An intuitive interpretation of $S_{k}$ is the point set that contains grid points inside and on the boundary of the bold line controlled by $\Gamma_{k}$, as shown in Figure LABEL:sub@fig:cft_2.
@@
  • If a new incumbent point $E_{i}$ is found with better SNR, the search region will be doubled as $\Gamma_{k+1}=2\Gamma_{k}$ (line 9), and the grid size will be empirically set as $\gamma_{k+1}=\min(\Gamma_{k},{\Gamma_{k}}^{2})$ (line 17), enabling the search in a broader space in the next iteration.
  • If there is no better point than the current $E_{k}$ on the current grids $G_{k}$, another searching stage will be performed only within the search region $S_{k}$ (line 10). Then, If a better point $E_{i}$ is found, $\Gamma_{k+1}$ and $\gamma_{k+1}$ are obtained as above (line 12), otherwise, the search region will be halved as $\Gamma_{k+1}=\Gamma_{k}/2$ (line 15) for a finer search with $\gamma_{k+1}=\min(\Gamma_{k},{\Gamma_{k}}^{2})$ (line 17).
@@
[…]
@@ III-D2 Pre-text Task Training and Fine-tuning
The pre-text task used for SSL should reveal certain inherent features in the radar-monitored vital sign and can be transferred to help the ECG recovery. According to our previous work [8], two major features of a good ECG recovery are accurate morphological features and accurate peak locations. The learning of morphological features fully relies on the non-linear mapping ability of the DL model by training with massive ECG data and is impossible to be pre-trained using radar signal only. However, the peaks in radar (heartbeats) and ECG signals (R peaks) are synchronized and have been extracted when assessing the radar signal SNR, as shown in Figure LABEL:sub@fig:target_sig.
It is natural to leverage the task for heartbeat detection as the pre-text task for SSL. Two categories of methods used for traditional heart rate estimation are based on periodicity and sparsity [2]. In this work, the duration of each segment is $4$ sec and may not reveal strong periodicity. Therefore, SSR will be used as the pre-text task in RFcardi and is defined as:
@@
where $h$ is the high-SNR radar signal, $\Phi$ means the observation matrix, $x$ is the sparse representation for heartbeats and $n$ represents the residual noise. The traditional SSR task can be seamlessly converted to a system identification problem by viewing $\Phi$ as a multi-channel adaptive filter, and the estimation of the filter coefficient is the same as training a CNN-based neural network (i.e., training CNN kernels) [30].
@@
$\mathcal{L}=\|x-\tilde{x}\|_{2}+\underbrace{\lambda_{s}\frac{\|x\|_{1}/\|x\|_{2}-1}{\sqrt{m}-1}}_{\text{sparse penalty}}$ (10)
where $m$ is the length of the signal, $x$ is the output from deep learning model, $\tilde{x}$ is the sparse ground truth with values for the radar peaks maintained (1st vibration in Figure LABEL:sub@fig:target_org) while other values are set to 0. The sparse penalty has a range of $0,\lambda_{s})$ with a smaller value indicating better sparsity [31].